OpenStudy (anonymous):

If f(x) = 4x – 1 and g(x) = x2 + 2x + 2, what is f(g(x))? can some explain. I keep getting the wrong answer... x2 + 8x – 2 x2 + 6x + 1 4x2 + 6x – 1 4x2 + 8x + 7

Parth (parthkohli):

Pre-note: \(f(g(x)) \ne f(x) \times g(x)\)

Parth (parthkohli):

\( \color{Black}{\Rightarrow f(g(x)) = f(x^2 + 2x + 2) = 4(x^2 + 2x + 2) - 1}\)

Parth (parthkohli):

Does that make it easier for you to complete the rest?

OpenStudy (anonymous):

f(x)=4x-1 here x is variable like f(t)=4t-1 Therefore f(g(x))=4(x^2+2x+2)-1 =4x^2+8x+7. ur ans seems to be right

OpenStudy (anonymous):

okay ty I wasn't sure.

Parth (parthkohli):

You're welcome!